# Design principles of the cytotoxic CD8+ T-cell response

### v1

@@ Abstract
Cytotoxic T lymphocytes eliminate infected or malignant cells, safeguarding surrounding tissues. Although experimental and systems-immunology studies have cataloged many molecular and cellular actors involved in an immune response, the design principles governing how the speed and magnitude of T-cell responses emerge from cellular decision-making remain elusive. Here, we recast a T-cell response as a feedback-controlled program, wherein the rates of activation, proliferation, differentiation and death are regulated through antigenic, pro- and anti-inflammatory cues. We demonstrate how the speed and magnitude of T-cell responses emerge from optimizing signal-feedback to protect against diverse infection settings. We recover an inherent trade-off: infection clearance at the cost of immunopathology. We show how this trade-off is encoded into the logic of T-cell responses by hierarchical sensitivity to different immune signals. This model explains experimentally observed patterns of T-cell effector expansion regulation in mice. Extending our model to immune-based T-cell therapies for cancer tumors, we identify a trade-off between the affinity for tumor antigens (“quality”) and the abundance (“quantity”) of infused T-cells necessary for effective treatment. Finally, we show how therapeutic efficacy can be improved by targeted genetic perturbations to T-cells. Our findings offer a unified control-logic for cytotoxic T-cell responses and point to specific regulatory programs that can be engineered for more robust T-cell therapies.
@@ I Introduction
We then ask how alternative immunological objectives give rise to distinct effector programs. First, we parameterize a broad class of feedback-controller designs, with varied signal sensitivities and integration rules. Next, by systematically sampling designs, we optimize for rapid pathogen clearance while minimizing immunopathology. We find that optimal designs lie on a Pareto front: further reductions in infection-induced harm entail increased immunopathology. Accounting for the imperative to separate infection—microbial or malignant—from transient inflammatory events (e.g., tissue damage) [28, 41, 17, 54], we show how a trade-off between pathogen-clearance and immunopathology (toxicity) may be encoded into the logic of T-cell responses. Optimizing these competing objectives yields effector programs consistent with experiments with vertebrate organisms. Finally, applying our framework to a minimal tumor-growth model, we demonstrate how modifying the endogenously encoded effector program can improve the effectiveness of T-cell cancer therapies, pointing to signal-processing adjustments—and thus candidate genetic engineering—that could enhance T-cell-based cancer immunotherapy.
@@
The immune system’s true performance objective, and how evolution has fine-tuned it, may ultimately be unknowable. Formulating and testing principled heuristics provides a baseline to gauge the performance of real immune systems in responding to infections. This perspective has yielded deep insights in other biological contexts, revealing the architectural logic of genetic and regulatory networks [23, 67, 42, 31, 6]. In immunology, high-level mathematical models have demonstrated how a finite lymphocyte repertoire can be distributed to best confront a diverse pathogenic landscape [59, 50, 11] and how immune memory should be allocated to counter future threats [51, 64, 15, 65]. By extending this perspective to the dynamics of CD8+ T-cell signal integration and response, our work provides a quantitative framework that links observed decision-making to operating principles and trade-offs.
@@
Figure 1: CD8+ T-cell response to infections. (A) The schematic shows how upon infection, susceptible cells $S$ become infected II at rate $b_{I}$. Infected cells provide an antigen signal $\sigma_{\text{Ag}}\propto I$, and die at rate $d_{I}$, which triggers an infection harm signal $\sigma_{\text{inf}}$. In response to infection, naïve CD8+ T-cells (with initial clone size of $N_{\text{lin.}}$) are activated $N\to N^{*}$, and can then differentiate into effector or memory fates. Effector cells can proliferate and kill infected and bystander cells, generating a response-harm signal $\sigma_{\text{res}}$, and finally become terminal or differentiate into memory. Consequently, T-cell transitions are modulated by the threes signals ($\sigma_{\text{Ag}},\sigma_{\text{inf}},\sigma_{\text{res}}$). Infection dynamics $(S,I)$ are modeled by ODEs coupled to a stochastic dynamics of the T-cell response. Memory transitions (gray arrows) are neglected in our model. (B) Representative trajectories of susceptible, infected, effector, and memory CD8+ T-cells are shown over time. (C) Each T-cell transition has a rate $r_{i}=r^{\text{max}}g_{i}(\sigma_{\text{Ag}},\sigma_{\text{inf}},\sigma_{\text{res}};\vec{\psi},\ell)$, modulated by a sigmoidal function with baseline $\ell$ and signal sensitivities $\vec{\psi}$ (eqs. 6, 7). The regulatory parameters define the design of the effector program $\vec{\theta}$ (eq. 8). (D) We assess design performance over diverse infections sampled from a broad range of reproduction numbers $R_{0}$ (eq. 2) and immunogenicities $\kappa$ (eq. 3). For each infection, we can find an optimal design $\vec{\theta}^{(\mathrm{min})}$ that minimizes total cell loss (harm). Harm decomposes into infection harm $H_{\text{inf.}}$ (cells killed by the pathogen plus infected cells eliminated by effectors), and toxicity $H_{\text{toxic.}}$ (healthy bystanders killed by effectors), highlighting an intrinsic trade-off (eq. 9). Simulation parameters in (B): $S_{\text{max}}=10^{7},N_{\text{lin.}}=100,K_{I}=10^{-3}S_{\text{max}},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},b_{I}=2.0\times 10^{-7}\text{day}^{-1}\penalty 10000\ \text{per cell},d_{I}=0.5\penalty 10000\ \text{day}^{-1},I_{0}=10^{-4}S_{\text{max}},\vec{\theta}=(\psi_{Ag}=2,\psi_{I}=2,\psi_{E}=-0.25,\ell_{N\to N^{*}}=-0.75,\ell_{N^{*}\to E}=-0.75,\ell_{E\to\emptyset}=0.75$).
@@
We model a population of healthy, susceptible cells $S$, initially consisting of $S_{\text{max}}$ identical cells. Pathogen exposure seeds an infected subpopulation $I(t)$, which grows via encounters with susceptible cells at rate $b_{I}$, and declines by infected-cell death at rate $d_{I}$, triggering an immune response (Fig. 1A). The dynamics of the susceptible and infected cells follow,
@@
where $d_{E}$ is the maximal clearance rate of infected/susceptible cells by effector T-cells $E$, and $K_{I}$ and $K_{S}$ are inverse affinity constants quantifying how well effector cells bind to and kill infected and susceptible cells, respectively [14, 25, 30]. The functional form of the last term in these equations can be motivated by considering competitive binding of T-cells to infected and healthy cells (SI Appendix, section 1).
@@
measuring the expected number of secondary infected cells generated by a single infected cell in an otherwise fully susceptible cell population. This definition mirrors the standard epidemiological reproduction number [55, 2]. We model the T-cell repertoire as initially consisting of a number of identical naïve T-cells, and assume $K_{S}=S_{\text{max}}\geq K_{I}$ so that infected cells are typically easier to recognize than susceptible cells. These inverse affinity parameters determine the immunogenicity of an infection, which we define as,
@@
This dimensionless quantity reflects how readily cognate T-cells recognize antigens presented by infected cells as “non-self” relative to those presented by healthy cells (self). We consider infection scenarios with broad ranges of immunogenicity ($\kappa$) and basic reproduction numbers ($R_{0}$) (SI Appendix, section 1).
@@
We simplify this picture by partitioning non-antigen cues into (i) infection-induced and (ii) response-induced harms, whose accumulated effects trigger infection and response harm-induced signals, $\sigma_{\text{inf}}(t)$ and $\sigma_{\text{res}}(t)$, respectively. Antigens presented by cells also serve as a signal; we take the infected-cell burden $I(t)$ as a proxy for the instantaneous antigen signal $\sigma_{\text{Ag}}(t)$. Together, expressed by their inverse affinity constants, $K_{I}$ for antigens and $K_{H}$ for non-antigens (harm), the three signals available to T-cells follow,
@@
The time-integrated harm-induced signals are exponentially discounted at rate $d_{H}$, reflecting the decay of cytokines in the absence of ongoing stimulation (SI Appendix, section 1).
@@
Naïve T-cells cycle through lymphoid tissues, becoming activated when they durably engage cognate peptide-MHC complexes on activated APCs [71, 24]. Activated T-cells can differentiate into an effector phenotype, which participates in infection clearance, or into a long-lived memory state (Fig. 1A). Here, we examine the course of CD8+ T-cells response during an infection, dividing the process into four canonical response modules (Figure. 1A):
@@
The parameter $\ell_{i}$ sets the baseline propensity of transition $i$: in the absence of external signals, the regulatory factor evaluates to $g_{0,i}=(1+\exp(-\ell_{i}))^{-1}$ (Figure. 1C). The weights $\{\psi_{i}^{Ag},\psi_{i}^{\text{inf}},\psi_{i}^{\text{res}}\}$ encode the sensitivity and polarity (positive / negative = up /down-regulation) of response to the antigen ($\sigma_{\text{Ag}}$) and harm signals ($\sigma_{\text{inf}}$ and $\sigma_{\text{res}}$). This formulation abstracts away biochemical details while implementing the fold-change response to stimuli observed in many biological circuits [31, 58, 3, 74, 4].
@@
Figure 2: Design parameters shape T-cell response and infection clearance. (A) Varying the design parameters influences effector trajectories, summarized by four features: activation time, expansion rate, peak effector magnitude, and contraction rate. (B) Trajectories are shown for the population size of effectors (top) and infected cells (bottom)—measured in units of maximum number of susceptible cells $S_{\text{max}}$—for designs that vary a single baseline transition rate $g_{0,i}$ (color) indicated atop each panel. All other design parameters are fixed at: fast activation, $g_{0,N\to N^{*}}=0.95$; fast differentiation, $g_{0,N^{*}\to E}=0.95$; slow contraction, $g_{0,E\to\emptyset}=0.05$; and zero signal feedback, $\psi_{\text{Ag}}=\psi_{\text{inf}}=\psi_{\text{res}}=0$. Trajectories reflecting the impact of varying signal sensitivities $\psi$’s are shown in SI Appendix, Fig. S1. (C) Sensitivities of trajectory (as in B) fold expansion, defined as the maximum effector count divided by the initial number of T-cells $N_{\text{lin.}}$, to changes in baseline transition rates (colors). Parameters are the same as Fig. 1, except for $K_{I}=10^{-2}S_{\text{max}}$.
@@
We systematically sample $\vec{\theta}$’s to assess performance of designs against different infections (Fig. 1D; SI Appendix, section 3). While $\vec{\theta}$ is broadly sampled, we hold the remaining biological rates and network parameters underlying the decision-making program in Fig. 1A within empirically supported ranges reported in SI Appendix, Table S1.
@@
where $H_{\text{inf.}}(\vec{\theta})$ aggregates infection-related harm (cells killed by the pathogen, and infected cells eliminated by effectors), $H_{\text{toxic.}}(\vec{\theta})$ captures immunopathology (healthy bystander cells killed by effectors), over the time window $T$ (SI Appendix, section 1 and 2, and Table S1). This single, interpretable metric balances clearance benefits against collateral damage, enabling principled comparison across designs and infections (Fig. 1D).
Figure 3: Toxicity-infection harm trade-off across T-cell regulatory designs. (A) Density plots of toxicity $H_{\text{toxic.}}$ versus infection harm $H_{\text{inf.}}$ (eq. 9)—measured in units of $S_{\text{max}}$—for $4.8\times 10^{6}$ T-cell regulatory designs, $\vec{\theta}$, uniformly grid-sampled across four infection scenarios (panels; columns: $\kappa=5.6$ (left), $\kappa=178$ (right); rows: $R_{0}=2.5$ (bottom), $R_{0}=5.0$ (top)). Designs are sampled uniformly across signal sensitivities $\psi\,\text{'s}\in[-3,3]$ and baseline rates $\ell\,\text{'s}\in[-3,3]$ (eq. 8). Infection scenarios are sampled by changing $K_{I}$ and $b_{I}$, while holding $S_{\max}$, $d_{I}$, and $K_{S}$ fixed. Black points denote no-feedback designs ($\vec{\psi}=0$), and the large magenta circle in each panel marks the design that minimizes total harm, $\vec{\theta}^{(\mathrm{min})}$. (B) For many different infection scenarios $(\kappa,R_{0})$, we identify an ensemble of near-optimal designs $\Theta(\kappa,R_{0})$, whose total harm is within $10^{-2}\times S_{\text{max}}$ of the minimum sampled total harm (as in the magenta point in A). By computing the ensemble-averaged design parameters $\langle\vec{\theta}\rangle_{\Theta(\kappa,R_{0})}$ for each infection scenario, and using these averages to cluster infections, we identify four infection “archetypes” (colors): I, II, III, IV (SI Appendix, section 3). (C) Toxicity vs. infection harm associated with optimal designs for each archetype is shown. Large dots represent cluster averages and small transparent dots represent individual infections, sampled in (B). (D) Bars indicate the ensemble-mean of design parameters, $\langle\vec{\theta}\rangle_{\Theta^{(i)}}$, computed by identifying the ensemble of designs $\Theta^{(i)}$} which minimize cluster-averaged total harm for archetype $i$ to within $10^{-2}\times S_{\text{max}}$ of the minimum sampled cluster-averaged total harm (similar to (B)). Error bars denote the standard deviation within this ensemble. All design variables are normalized by their respective bounds ($\psi^{\max}=3$, $\ell^{\max}=3$). Unless otherwise noted, simulation parameters: $S_{\text{max}}=10^{7},N_{\text{lin.}}=10^{-5}S_{\text{max}},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},d_{I}=0.5\penalty 10000\ \text{day}^{-1},I_{0}=10^{-4}S_{\text{max}}.$

### v2

@@ Abstract
Cytotoxic T lymphocytes eliminate infected or malignant cells, safeguarding surrounding tissues. Although experimental and systems-immunology studies have cataloged many molecular and cellular actors involved in an immune response, the design principles governing how the speed and magnitude of T-cell responses emerge from cellular decision-making remain elusive. Here, we recast the T-cell response as a feedback-controlled program, wherein the rates of activation, proliferation, differentiation and death are regulated through antigenic, pro- and anti-inflammatory cues. By exploring a broad class of feedback-controller designs as potential immune programs, we demonstrate how the speed and magnitude of T-cell responses emerge from optimizing signal-feedback to protect against diverse infection settings. We recover an inherent trade-off: infection clearance at the cost of immunopathology. We show how this trade-off is encoded into the logic of T-cell responses by hierarchical sensitivity to different immune signals. Notably, we find that designs that balance harm from acute infections and autoimmunity produce immune responses consistent with experimentally observed patterns of T-cell effector expansion in mice. Extending our model to immune-based T-cell therapies for cancer tumors, we identify a trade-off between the affinity for tumor antigens (“quality”) and the abundance (“quantity”) of infused T-cells necessary for effective treatment. Finally, we show how therapeutic efficacy can be improved by targeted genetic perturbations to T-cells. Our findings offer a unified control-logic for cytotoxic T-cell responses and point to specific regulatory programs that can be engineered for more robust T-cell therapies.
@@ I Introduction
We then ask how alternative immunological objectives give rise to distinct effector programs. First, we parameterize a broad class of feedback-controller designs, with varied signal sensitivities and integration rules. Next, by systematically sampling designs, we optimize for rapid pathogen clearance while minimizing immunopathology. We find that optimal designs lie on a Pareto front: further reductions in infection-induced harm entail increased immunopathology. Accounting for the imperative to separate infection—microbial or malignant—from transient inflammatory events (e.g., tissue damage) [25, 26, 27, 28, 29], we show how a trade-off between pathogen-clearance and immunopathology (toxicity) may be encoded into the logic of T-cell responses. Optimizing these competing objectives yields effector programs consistent with experiments with vertebrate organisms. Finally, applying our framework to a minimal tumor-growth model, we demonstrate how modifying the endogenously encoded effector program can improve the effectiveness of T-cell cancer therapies, pointing to signal-processing adjustments—and thus candidate genetic engineering—that could enhance T-cell-based cancer immunotherapy.
@@
The immune system’s true performance objective, and how evolution has fine-tuned it, may ultimately be unknowable. Formulating and testing principled heuristics provides a baseline to gauge the performance of real immune systems in responding to infections. This perspective has yielded deep insights in other biological contexts, revealing the architectural logic of genetic and regulatory networks [30, 31, 32, 33, 34, 35]. In immunology, high-level mathematical models have demonstrated how a finite lymphocyte repertoire can be distributed to best confront a diverse pathogenic landscape [36, 37, 38] and how immune memory should be allocated to counter future threats [39, 40, 41, 42, 43]. By extending this perspective to the dynamics of CD8+ T-cell signal integration and response, our work provides a quantitative framework that links observed decision-making to operating principles and trade-offs.
@@
Figure 1: CD8+ T-cell response to infections. (A) The schematic shows how upon infection, susceptible cells $S$ become infected II at rate $b_{I}$. Infected cells provide an antigen signal $\sigma_{\text{Ag}}\propto I$, and die at rate $d_{I}$, which triggers an infection harm signal $\sigma_{\text{inf}}$. In response to infection, naïve CD8+ T-cells (with initial clone size of $N_{\text{lin.}}$) are activated $N\to N^{*}$, and can then differentiate into effector or memory fates (Fig. S1). Effector cells can proliferate and kill infected and bystander cells, generating a response-harm signal $\sigma_{\text{res}}$, and finally become terminal or differentiate into memory. Consequently, T-cell transitions are modulated by the threes signals ($\sigma_{\text{Ag}},\sigma_{\text{inf}},\sigma_{\text{res}}$). Infection dynamics $(S,I)$ are modeled by ODEs coupled to a stochastic dynamics of the T-cell response. Memory transitions (gray arrows) are neglected in our model. (B) Representative trajectories of susceptible, infected, effector, and memory CD8+ T-cells are shown over time. (C) Each T-cell transition has a rate $r_{i}=r^{\text{max}}g_{i}(\sigma_{\text{Ag}},\sigma_{\text{inf}},\sigma_{\text{res}};\vec{\psi},\ell)$, modulated by a sigmoidal function with baseline $\ell$ and signal sensitivities $\vec{\psi}$ (eqs. 6, 7). The regulatory parameters define the design of the effector program $\vec{\theta}$ (eq. 8). (D) We assess design performance over diverse infections sampled from a broad range of reproduction numbers $R_{0}$ (eq. 2) and immunogenicities $\kappa$ (eq. 3). For each infection, we can find an optimal design $\vec{\theta}^{(\mathrm{min})}$ that minimizes total cell loss (harm). Harm decomposes into infection harm $H_{\text{inf.}}$ (cells killed by the pathogen plus infected cells eliminated by effectors), and toxicity $H_{\text{toxic.}}$ (healthy bystanders killed by effectors), highlighting an intrinsic trade-off (eq. 9). Simulation parameters in (B): $S_{\text{max}}=10^{7},N_{\text{lin.}}=100,K_{I}=10^{-3}S_{\text{max}},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},b_{I}=2.0\times 10^{-7}\text{day}^{-1}\penalty 10000\ \text{per cell},d_{I}=0.5\penalty 10000\ \text{day}^{-1},I_{0}=10^{-4}S_{\text{max}},\vec{\theta}=(\psi_{Ag}=2,\psi_{I}=2,\psi_{E}=-0.25,\ell_{N\to N^{*}}=-0.75,\ell_{N^{*}\to E}=-0.75,\ell_{E\to\emptyset}=0.75$).
@@
We model a population of healthy, susceptible cells $S$, initially consisting of $S_{\text{max}}$ identical cells. Pathogen exposure seeds an infected subpopulation $I(t)$, which grows via encounters with susceptible cells at rate $b_{I}$, and declines by infected-cell death at rate $d_{I}$, triggering an immune response (Figs. 1A, S1). The dynamics of the susceptible and infected cells follow,
@@
where $d_{E}$ is the maximal clearance rate of infected/susceptible cells by effector T-cells $E$, and $K_{I}$ and $K_{S}$ are recognition thresholds, quantifying how readily effector cells can recognize infected and susceptible cells, respectively, and consequently kill them; a larger $K_{I}$ or $K_{S}$ implies more effector cells are needed for recognition [44, 45, 46]. The functional form of the last term in these equations can be motivated by considering competitive binding of T-cells to infected and healthy cells (SI Appendix).
@@
measuring the expected number of secondary infected cells generated by a single infected cell in an otherwise fully susceptible cell population. This definition mirrors the standard epidemiological reproduction number [47, 48]. We model the T-cell repertoire as initially consisting of a number of identical naïve T-cells, and assume $K_{S}=S_{\text{max}}\geq K_{I}$ so that infected cells are typically easier to recognize than susceptible cells. These recognition thresholds determine the immunogenicity of an infection, which we define as,
@@
This dimensionless quantity reflects how readily cognate T-cells recognize antigens presented by infected cells as “non-self” relative to those presented by healthy cells (self). We consider infection scenarios with broad ranges of immunogenicity ($\kappa$) and basic reproduction numbers ($R_{0}$) (SI Appendix).
@@
We simplify this picture by partitioning non-antigen cues into (i) infection-induced and (ii) response-induced harms, whose accumulated effects trigger infection and response harm-induced signals, $\sigma_{\text{inf}}(t)$ and $\sigma_{\text{res}}(t)$, respectively. Antigens presented by cells also serve as a signal; we take the infected-cell burden $I(t)$ as a proxy for the instantaneous antigen signal $\sigma_{\text{Ag}}(t)$. Together, expressed by their recognition thresholds, $K_{I}$ for antigens and $K_{H}$ for non-antigens (harm), the three signals available to T-cells follow,
@@
The time-integrated harm-induced signals are exponentially discounted at rate $d_{H}$, reflecting the decay of cytokines in the absence of ongoing stimulation (SI Appendix).
@@
Naïve T-cells cycle through lymphoid tissues, becoming activated when they durably engage cognate peptide-MHC complexes on activated APCs [49, 50]. Activated T-cells can differentiate into an effector phenotype, which participates in infection clearance, or into a long-lived memory state (Figs. 1A, S1). Here, we examine the course of CD8+ T-cells response during an infection, dividing the process into four canonical response modules (Fig. 1A):
@@
The parameter $\ell_{i}$ sets the baseline propensity of transition $i$: in the absence of external signals, the regulatory factor evaluates to $g_{0,i}=(1+\exp(-\ell_{i}))^{-1}$ (Fig. 1C). The weights $\{\psi_{i}^{Ag},\psi_{i}^{\text{inf}},\psi_{i}^{\text{res}}\}$ encode the sensitivity and polarity (positive / negative = up /down-regulation) of response to the antigen ($\sigma_{\text{Ag}}$) and harm signals ($\sigma_{\text{inf}}$ and $\sigma_{\text{res}}$). This formulation abstracts away biochemical details while implementing the fold-change response to stimuli observed in many biological circuits [33, 54, 55, 56, 57].
@@
Figure 2: Design parameters shape T-cell response and infection clearance. (A) Varying the design parameters influences effector trajectories. (B) Trajectories are shown for the population size of effectors (top) and infected cells (bottom), measured in units of maximum number of susceptible cells $S_{\text{max}}$, for designs that vary a single baseline transition rate $g_{0,i}$ (color) indicated atop each panel. All other design parameters are fixed at: fast activation, $g_{0,N\to N^{*}}=0.95$; fast differentiation, $g_{0,N^{*}\to E}=0.95$; slow contraction, $g_{0,E\to\emptyset}=0.05$; and zero signal feedback, $\psi_{\text{Ag}}=\psi_{\text{inf}}=\psi_{\text{res}}=0$. Trajectories reflecting the impact of varying signal sensitivities $\psi$’s are shown in SI Appendix, Fig. S2. (C) Peak effector count, measured in units of maximum number of susceptible cells $S_{\text{max}}$, is shown by varying each of the three baseline transition rates (colors), while keeping the rest of the parameters fixed (similar to (B)). Parameters are the same as Fig. 1, except for $K_{I}=10^{-2}S_{\text{max}}$.
@@
We systematically sample $\vec{\theta}$’s to assess performance of designs against different infections (Fig. 1D; SI Appendix). While $\vec{\theta}$ is broadly sampled, we hold the remaining biological rates and design parameters underlying the decision-making program in Fig. 1A within empirically supported ranges reported in SI Appendix, Table S1.
@@
where $H_{\text{inf.}}(\vec{\theta})$ aggregates infection-related harm (cells killed by the pathogen, and infected cells eliminated by effectors), $H_{\text{toxic.}}(\vec{\theta})$ captures immunopathology (healthy bystander cells killed by effectors), over the time window $T$ (SI Appendix, Table S1). This single, interpretable metric balances clearance benefits against collateral damage, enabling principled comparison across designs and infections (Fig. 1D).
Figure 3: Toxicity-infection harm trade-off across T-cell regulatory designs. (A) Density plots of toxicity $H_{\text{toxic.}}$ versus infection harm $H_{\text{inf.}}$ (eq. 9)—measured in units of $S_{\text{max}}$—for $4.8\times 10^{6}$ T-cell regulatory designs, $\vec{\theta}$, uniformly grid-sampled across four infection scenarios (panels; columns: $\kappa=5.6$ (left), $\kappa=178$ (right); rows: $R_{0}=2.5$ (bottom), $R_{0}=5.0$ (top)). Designs are sampled uniformly across signal sensitivities $\psi\,\text{'s}\in[-3,3]$ and baseline rates $\ell\,\text{'s}\in[-3,3]$ (eq. 8). Infection scenarios are sampled by changing $K_{I}$ and $b_{I}$, while holding $S_{\max}$, $d_{I}$, and $K_{S}$ fixed. Black points denote no-feedback designs ($\vec{\psi}=0$), and the large magenta circle in each panel marks the design that minimizes total harm, $\vec{\theta}^{(\mathrm{min})}$. (B) For many different infection scenarios $(\kappa,R_{0})$, we identify an ensemble of near-optimal designs $\Theta(\kappa,R_{0})$, whose total harm is within $10^{-2}\times S_{\text{max}}$ of the minimum sampled total harm (as in the magenta point in A). By computing the ensemble-averaged design parameters $\langle\vec{\theta}\rangle_{\Theta(\kappa,R_{0})}$ for each infection scenario, and using these averages to cluster infections, we identify four infection “archetypes” (colors): I–IV, as indicated (SI Appendix). (C) Toxicity vs. infection harm associated with optimal designs for each archetype is shown. Large dots represent cluster averages and small transparent dots represent individual infections, sampled in (B). (D) Bars indicate the ensemble-mean of design parameters, $\langle\vec{\theta}\rangle_{\Theta^{(i)}}$, computed by identifying the ensemble of designs $\Theta^{(i)}$} which minimize cluster-averaged total harm for archetype $i$ to within $10^{-2}\times S_{\text{max}}$ of the minimum sampled cluster-averaged total harm (similar to (B)). Error bars denote the standard deviation within this ensemble. All design variables are normalized by their respective bounds ($\psi^{\max}=3$, $\ell^{\max}=3$). Unless otherwise noted, simulation parameters: $S_{\text{max}}=10^{7},N_{\text{lin.}}=10^{-5}S_{\text{max}},K_{S}=S_{\text{max}},K_{H}=10^{-3}S_{\text{max}},d_{I}=0.5\penalty 10000\ \text{day}^{-1},I_{0}=10^{-4}S_{\text{max}}.$